Line A fills tank in 3 hours. Line B fills tank in 4 hours. Line C empties tank in 5 hours. With all lines open how long until tank is 3/5 full
Answer by ewatrrr(24785)


Hi,
Let x represent the time it would take to fill tank for all three working together.
Question states; Per Hour the Equalizer:
1/3 + 1/4 - 1/5 = 1/x Multiplying each term by 60x so as all denominators = 1
20x + 15x - 12x = 60
23x = 60
x = 60/23 hr
With all lines open how long until tank is 3/5 full
3/5 * 60/23 = 36/23 = 1.57 hrs, tank will be 3/5 full